« Project Status: 9 January 2008 0100UTC (IOP#5 Update) | Main | Project Status: 10 January 2008 2030UTC »

Project Status: 9 January 2008 2030UTC

IOP#5 Comes to an End

Weather Status: The fast moving storm of yesterday barreled though the ARB from late morning 8 Jan 17UTC (09L) to early this morning 9 Jan 10UTC (02L). Precipitation amounts around the area were 1 to 2 inches of liquid, about 10 inches of snow at BLU. BLU remained in snow the entire period, getting a dose of ice at one point when supercooled cloud put some ice on the dish. Snow to liquid ratios were about 6 to1 becoming wetter in the latter part of the event (5 to 1). The freezing level that started out at about 3000ft seemed to rise through the event as evidenced by the increase from 3000 to 5000 ft at Chico. By the end of the event the temperature at BLU was -1 to 0C (30-31F)so it was likely the melting layer was just below. The next wave due to arrive tonight and Thursday continues to look weak on the numerical guidance. GFS and NAM estimates are 0.5 inches or so. There is an upper level jet over the ARB so increased amounts in bands might be expected. A weak trough moving north of the area on Saturday could bring scattered showers to the central Sierras, but nothing warranting operations. Beyond that GFS and ECMWF ensemble members are cutting off a low in the east Pacific that may drift east. This will bear watching next week. All in all pretty quiet for the next 10 days.

Summary: No system of interest for the next 5 days.

Observational Status: Sondes successfully released at SHS, RNO, and OAK from 8 Jan 12UTC (04L) to 9 Jan 03 UTC (19L) at 3-hour intervals; Radar was operational from 8 Jan 1530UTC (0730L) to 9 Jan 05UTC (21L). Radar crew departed the site at 8/22L and are returning to Boulder today. BLU surface site was down for the IOP but came up today; Lee Vining surface site down. Cazadero profiler down.

Decision Summary: No operations planned for the next 5 days.

Product Status: NWP high resolution guidance did well on QPF, but had the precip into the ARB a bit early. National NWP guidance was very helpful with good QPF and timing. ALPS Workstation news: GSD personnel finishing up work on software. Workstation will be shipped to MTY and RNO Friday (11Jan); ALPS software for SAC will be tarred and FTP'd Friday.

McGinley, GSD

TrackBack

TrackBack URL for this entry:
http://www.etl.noaa.gov/cgi-bin/mt/mt-tb.cgi/716